OpenStudy (anonymous):

cilia cycles 4.2 km every morning. how many feet are in 4.2 km, given that 1 mile = 1.609 km and 1 mile = 5,280 feet?

OpenStudy (anonymous):

Alright first you'd want to convert the 4.2 km to miles.

OpenStudy (anonymous):

4.2km/1.609km/mile=2.610 miles. Ok, so we know there are 2.610 miles in 4.2km

OpenStudy (anonymous):

Now since 1 mile=5,280 feet, you have to multiply 2.610 miles with 5280 feet

OpenStudy (anonymous):

13780.8 feet

OpenStudy (anonymous):

thats not one of the answer choices

OpenStudy (anonymous):

ok. type all of the possible choices.

OpenStudy (mathstudent55):

\(\large 4.2 ~km \times \dfrac{1~mile}{1.609~km} \times \dfrac{5280~ft}{1~mile} \)

OpenStudy (anonymous):

Good work @mathstudent55

OpenStudy (mathstudent55):

4.2 km = 13,782 ft

OpenStudy (mathstudent55):

@Robert136 Thanks, you too.

OpenStudy (anonymous):

13,782.47 22,176.00 35,681.18 44,352.00